3. How would the following changes in price affect total revenue? That is, would total revenue increase, decrease, or remain unchanged?

a. Price falls and demand is inelastic.

b. Price rises and demand is elastic.


4. The price elasticity of demand for beef is about 0.60. Other things equal, this means that a 20 percent increase in the price of beef will cause the quantity of beef demanded to____.


5. Carlsberg recorded an increase of 33% of net profit for the 1st quarter of 2016, despite of increasing taxes.” Why? Could you provide a reasonable explanation using both economic and business knowledge..



1
Expert's answer
2021-02-03T16:10:54-0500

Solution:

3.). a. Price falls and demand is inelastic.

If the demand is inelastic at the original quantity level, a fall in price will cause a slight increase in quantity sold and so the total revenue will decrease.

b. Price rises and demand is elastic.

If the demand is elastic at the original quantity level, an increase in price will be more than offset by a larger fall in quantity resulting in a fall in revenue.

 

4.). The quantity of beef demanded will decrease by approximately 12 percent.


The price elasticity of demand = "\frac{\%\;change \;in\;quantity\;demanded}{\%\;change \;in\;price}"


"0.60 = \frac{\%\;change \;in\;Qd}{0.20}"


"\%\;change \;in\;Qd = 0.60\times 0.20 = 0.12\times 100 = 12\%"

 

5.). The reason why Carlsberg recorded an increased net profit percentage for the 1st quarter of 2016 is because of an increase in total revenue and a decrease in operating expenses resulting in a higher gross margin. The reason for the increase in revenue is because Carlsberg increased the prices of its products and while maintaining its sales. This countered the effect of increased taxes resulting in an increased profit for the 1st quarter of 2016.